### v2

# DHP: Discrete Hierarchical Planning for HRL Agents

Abstract
Hierarchical Reinforcement Learning (HRL) agents often struggle with long-horizon visual planning due to their reliance on error-prone distance metrics. We propose Discrete Hierarchical Planning (DHP), a method that replaces continuous distance estimates with discrete reachability checks to evaluate subgoal feasibility. DHP recursively constructs tree-structured plans by decomposing long-term goals into sequences of simpler subtasks, using a novel advantage estimation strategy that inherently rewards shorter plans and generalizes beyond training depths. In addition, to address the data efficiency challenge, we introduce an exploration strategy that generates targeted training examples for the planning modules without needing expert data. Experiments in 25-room navigation environments demonstrate $100\%$ success rate (vs $82\%$ baseline) and $73$-step average episode length (vs $158$-step baseline). The method also generalizes to momentum-based control tasks and requires only $\log N$ steps for replanning. Theoretical analysis and ablations validate our design choices.
1 Introduction
Hierarchical planning enables agents to solve complex tasks through recursive decomposition [8], but existing approaches face fundamental limitations in long-horizon visual domains. While methods using temporal distance metrics [27, 2] or graph search [10] have shown promise, their reliance on precise continuous distance estimation creates two key challenges:
  • Coupled Learning Dynamics: Distance estimates depend on the current policy’s quality, where suboptimal policies produce misleading distances and practical implementations may require arbitrary distance cutoffs [2, 10].
  • Exploratory Objective: The intrinsic rewards used for training explorers may not align with the planning objective leading to inaccurate distance measures by design (Fig. 10).
We address these issues by reformulating hierarchical planning through discrete reachability – a paradigm shift from "How far?" to "Can I get there?". Our method (DHP) evaluates plan feasibility through binary reachability checks rather than continuous distance minimization. This approach builds on two insights: local state transitions are easier to model than global distance metrics, and reachability naturally handles disconnected states through $0/1$ signaling.
Our key contributions:
  • A reachability-based discrete reward scheme for planning (Sec. 2.3.2, Fig. 8(d))
  • An advantage estimation algorithm for tree-structured trajectories (Sec. 2.3.3, Fig. 8(b)).
  • A memory conditioned exploration strategy outperforming expert data (Sec. 2.4, Figs. 8(c),11).
  • An extensive ablation study that measures the contribution of each module (Sec. 3.2).
Empirical results confirm these design choices. On the 25-room navigation benchmark, DHP achieves perfect success rates ($100\%$ vs $82\%$) with 2.1x shorter trajectories than state-of-the-art (Sec. 3.1). Our ablation studies demonstrate that both the reachability paradigm and the novel advantage estimation contribute significantly to these gains (Section 3.2). We also test our method in a momentum-based RoboYoga environment for generalizability. The code is available at: <URL redacted>.
2 Discrete Hierarchical Planning (DHP)
In the context of Markovian Decision Processes (MDP), a Reinforcement Learning task can be imagined as an agent transitioning through states $s_{t}$ using actions $a_{t}$. Our task is to find the shortest path between any two given states $(s_{t},s_{g})$. To do this, we first use an explorer to collect a dataset of useful trajectories possible within the environment. Then we learn a planning policy that learns to predict subgoals that decompose the initial task into two simpler sub-tasks. The recursive application of the policy breaks the subtasks further till subtasks directly manageable by the worker are found.
2.1 Agent Architecture
((a)) Agent architecture
((b)) GCSR module
((c)) Subgoal prediction using GCSR decoder
Figure 1: Illustrations for different module architectures. (a) Overall planning agent architecture. The world model predicts the state $s_{t}$, the planner takes the current and goal states $(s_{t},s_{g})$ to output a latent variable $z$, the GCSR Decoder is then used to predict a subgoal $s_{i}$. Then the subgoal is used as a goal to predict another subgoal. This continues recursively till a reachable subgoal $s_{wg}$ is found, which is then passed to the worker. (b) The GCSR module is a conditional VAE that consists of an encoder and a decoder optimized to predict midway states, given the initial and final states. (c) The planning policy uses the GCSR decoder to predict subgoals.
We use the base architecture from the Director [14], which consists of three modules: world-model, worker, and manager. World model is implemented using the Recurrent State Space Module (RSSM) [13] that learns state representations $s_{t}$ using a sequence of image observations $o_{t}$. The worker policy $\pi_{W}$ learns to output environmental actions $a_{t}$ to reach nearby states $s_{wg}$. The manager $\pi_{M}$ is a higher-level policy that learns to output desirable sub-goal states $s_{wg}$ for the worker (updated every $K$ steps) in the context of an external task and the exploratory objective jointly. We use our explorer policy $\pi_{E}$ and the planning policy $\pi_{P}$ as the manager. Figure 1(a) shows the agent architecture during inference using $\pi_{P}$. The worker and the world-model are trained using the default objectives from the Director (for more details see Appendix B). First, we describe our planning policy and then derive an exploratory strategy fit for it.
2.2 Planning Policy
The Planning policy $\pi_{P}$ is a goal-conditioned policy that takes the current and goal state as inputs to yield a subgoal. However, predicting directly in the continuous state space leads to the problem of high-dimensional continuous control [14, 27]. To reduce the search space for the planning policy, we train a discrete Conditional Variational AutoEncoder (CVAE) that learns to predict midway states $s_{t+q/2}$ given an initial and a final state $(s_{t},s_{t+q})$. We call this module Goal-Conditioned State Recall (GCSR). It consists of an Encoder and a Decoder (Fig. 1(b)). The Encoder takes the initial, midway, and final states $(s_{t},s_{t+q/2},s_{t+q})$ as input to output a distribution over a latent variable $\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})$. The Decoder uses the states $(s_{t},s_{t+q})$ and a sample from the latent distribution $z\sim\text{Enc}_{G}(z)$ to predict the midway state $\text{Dec}_{G}(s_{t},s_{t+q},z)\rightarrow\hat{s}_{t+q/2}$. The module is optimized using the explorer collected data to minimize the ELBO objective (Eq. 1). A mixture of categoricals $(4\times 4)$ as the prior latent distribution $p_{G}(z)$ for all our experiments. Triplets at multiple temporal resolutions $q\in Q$ are extracted, allowing the agent to function at all temporal resolutions. For minimum overlap between resolutions, we use exponentially increasing temporal resolutions, $Q=\{2K,4K,8K,...\}$.
$\begin{split}\mathcal{L}(\text{Enc}_{G},\text{Dec}_{G})=\sum_{q\in Q}\left\|s_{t+q/2}-\text{Dec}_{G}(s_{t},s_{t+q},z)\right\|^{2}+\beta\text{KL}[\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})\parallel p_{G}(z)]\\ \text{where}\quad z\sim\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})\end{split}$ (1)
The planning policy predicts in the learned latent space, which are expanded into sub-goals using the GCSR decoder as: $\text{Dec}_{G}(s_{t},s_{g},z)$ where $z\sim\pi_{P}(z|s_{t},s_{g})$ (Fig. 1(c)). Note that the word discrete in our title refers to our plan evaluation method (Sec. 2.3.2), not the action space of the planning policy.
((a)) Unrolled subtask tree during Training
((b)) Planning during Inference
Figure 2: The figure illustrates the plan unrolling process during the training and inference phases. (a) During Training, the initial subtask $(s_{t},s_{g})$ is recursively decomposed into smaller tasks by midway subgoal prediction to generate a subtask tree. The lowest level nodes represent the simplest decomposition of the initial task as: $(s_{t},s_{g})\rightarrow(s_{t},s_{3},s_{1},s_{4},s_{0},s_{5},s_{2},s_{6},s_{g})$. (b) During Inference, only the first branch of the tree is unrolled. Here, the agent is tasked to get to $s_{g}$. So it first divides the task $(s_{t},s_{g})$ into two chunks by inserting $s_{0}$. Then it proceeds to divide the subtask $(s_{t},s_{0})$ by inserting $s_{1}$ and ignoring the second part $(s_{0},s_{g})$. The recursive division continues till the first subgoal reachable in $K$ steps is found. This results in a stack of subgoals shown at the bottom.
2.3 Planning Policy Optimization
The planning policy is optimized as a Soft-Actor-Critic (SAC) [12] in three steps: construct plans between random state pairs (Sec. 2.3.1), plan evaluation using discrete rewards and our novel advantage estimation method (Sec. 2.3.2,2.3.3), and policy updates using policy gradients (Sec. 2.3.4).
2.3.1 Plan Unrolling
Given the initial and final states $(s_{t},s_{g})$, a subgoal generation method predicts an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator breaks the initial task, leading to a tree of subtasks $\tau$. The root node $n_{0}$ represents the original task $(s_{t},s_{g})$, and each remaining node $n_{i}$ in the tree $\tau$ represents a sub-task. At each node $n_{i}$, the policy predicts a subgoal as: $s_{i}=\text{Dec}_{G}(n_{i,0},n_{i,1},z)$ where $z\sim\pi_{P}(z_{i}|n_{i})$. The preorder traversal of the subtask tree of depth $D$ can be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. Figure 2(a) shows an example unrolled tree. The lowest-level nodes show the smallest decompositions of the task under the current planning policy $\pi_{P}$.
Inference: Unlike traditional methods for hierarchical planning (eg, cross-entropic methods (CEM) [29, 22]), which require unrolling multiple full trees followed by evaluation at runtime, our method does not require full tree expansion. A learned policy always predicts the best estimated subgoal by default. This also means that we do not need to unroll the second subtask further at each node because only the first reachable subgoal is required (Fig. 2(b)).
2.3.2 Discrete Rewards Scheme
We want to encourage trees that end in subtasks manageable by the worker. A subtask is worker-manageable if the node goal $n_{i,1}$ is reachable by the worker from the node initial state $n_{i,0}$. Since learning modules inside other learned modules can compound errors, we use a more straightforward method to check reachability. We simulate the worker for $K$ steps using RSSM imagination, initialized at $n_{i,0}$ with goal $n_{i,1}$. If the cosine_max similarity measure between the worker’s final state $s_{t,i}$ and the assigned goal state is above a threshold $\Delta_{R}$, the node is marked as terminal (Eq. 2). The terminal nodes do not need further expansion, which is different from the word’s usual meaning in the context of trees. And the lowest layer nodes in a finite depth unrolled tree are called truncated nodes. Computing the terminal array $T_{i}$ allows supporting imperfect trees with branches terminating at different depths. A plan is considered valid if all its branches end in terminal nodes. The policy is rewarded $1$ at terminal nodes and 00 otherwise (Eq. 3). A discrete reward scheme allows for optimization that increases the likelihood of valid plans over distance-based approaches that gradually optimize the policy to reduce path length.
$T_{i}=$ $T_{(i-1)/2}\lor\texttt{cosine\_max}(s_{t,i},n_{i,1})>\Delta_{R}$ (2)
$R_{i}=$$\begin{cases}1,&\text{if }T_{i}==True\\ 0,&\text{otherwise}\end{cases}$ (3)
2.3.3 Return Estimation for Trees
Taking inspiration from the standard discounted return estimation for linear trajectories (sequence of states) [31], we propose an approach for tree trajectories. Returns for a linear trajectory are computed as the reward received at the next step and discounted rewards thereafter, $G_{t}=R_{t+1}+\gamma G_{t+1}$. Similarly, the return estimation for trees is the minimum discounted return from the child nodes. Given a tree trajectory $\tau$, the Monte-Carlo return, the $1$-step return, and the lambda return for each non-terminal node as can be written as:
$G_{i}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma G_{2i+1},R_{2i+2}+\gamma G_{2i+2})$ (4)
$G_{i}^{0}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma v_{P}(n_{2i+1}),R_{2i+2}+\gamma v_{P}(s_{2i+2}))$ (5)
$G_{i}^{\lambda}=(1-T_{i})\cdot\min(R_{2i+1}+\gamma((1-\lambda)v_{P}(n_{2i+1})+\lambda G_{2i+1}^{\lambda}),R_{2i+2}+\gamma((1-\lambda)v_{P}(s_{2i+2})+\lambda G_{2i+2}^{\lambda}))$ (6)
All branches should end in terminal nodes to score a high return with the above formulation. Also, since the discount factor diminishes the return with each additional depth, the agent can score higher when the constructed tree is shallow (less maximum depth). This characteristic is similar to linear trajectories, where the return is higher for shorter paths to the goal [33].
Fig. 3 illustrates example return evaluations for Monte-Carlo returns and $n$-step truncated returns that use value estimates to replace rewards at the non-terminal truncated nodes. The $n$-step returns allow for generalization beyond the maximum unrolled depth $D$ (Sec. A.2.2). Thus, the tree can be unrolled for higher depths $D_{I}$ during inference. We show that the Bellman operators for the above returns are contractions and repeated applications cause the value function $v^{\pi}_{P}$ to approach a stationary $v^{*}$ (Sec. A.4). We explore how the return penalizes maximum tree depth and encourages balanced trees in Sec. A.2.3, implying that the policy breaks tasks roughly halfway.
((a)) Monte-Carlo return
((b)) $n$-step return
Figure 3: Example return estimations for an imperfect tree (node indices at top-left). The dash-bordered cells indicate terminal nodes where the policy receives a $1$ reward and the branch terminates. While one of the branches terminates early $(i=4)$, one does not for the unrolled depth $(i=11)$. (a) Since we compute the return as the $\min$ of the child nodes, the Monte-Carlo return at the root node is 00 in this case. However, there is still a positive learning signal induced at nodes $(i=1,3,6)$. (b) Using the critic as a baseline for computing $n$-step returns. The $n$-step returns allow bootstrapping by substituting the reward with value estimates $v$ at the truncated node $(i=11)$. This induces a learning signal at the root node even if the plan is incomplete for the unrolled depth.
2.3.4 Policy Gradients
Using tree return estimates (we use $n$-step lambda returns in all cases), we derive the policy gradients for the planning policy as (Eq. 7, proof in Sec. A.1). We show that if the function $G^{\lambda}_{i}$ is replaced by a function independent of the policy actions, the expectation reduces to 00, implying that we can use the value function as a baseline for variance reduction (Th. A.2). With policy gradients and an entropy term, to encourage random exploration before convergence, we construct the loss function for the actor $\pi_{P}$ and critic $v_{P}$ as (sum over all nodes except the terminal and truncated nodes):
$\nabla_{\pi_{P}}J(\pi_{P})=\mathbb{E}_{\tau}\sum_{i=0}^{2^{D}-2}G_{i}^{\lambda}(\tau)\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})$ (7)
$\mathcal{L}(\pi_{P})=-\mathbb{E}_{\tau\sim\pi_{P}}\sum_{i=0}^{2^{D}-2}(1-T_{i})\cdot(G^{\lambda}_{i}-v_{P}(n_{i}))\log\pi_{P}(z_{i}|n_{i})+\eta\mathbb{H}[\pi_{P}(z_{i}|n_{i})]$ (8)
$\mathcal{L}(v_{P})=\mathbb{E}_{\tau\sim\pi_{P}}\sum^{2^{D}-2}_{i=0}(1-T_{i})\cdot(v_{P}(n_{i})-G^{\lambda}_{i})^{2}$ (9)
2.4 Explorer
During exploration, an exploration policy $\pi_{E}$ is used as a manager to drive the worker behavior. For the same problem of continuous control, the Explorer predicts goals in a discrete latent space learned using a VAE. As the predicted states do not need to be conditioned on other states like the planner, the Explorer VAE learns state representations unconditionally (similar to Director). The Explorer VAE consists of an encoder and a decoder $(\text{Enc}_{U},\text{Dec}_{U})$. The encoder predicts latent distributions using state representations: $\text{Enc}_{U}(z|s_{t})$, and the decoder tries to reconstruct the states using the samples from the latent distribution: $\text{Dec}_{U}(z)$ (Fig. 5(a)). As the prediction space is unconstrained, we use a larger latent size ($8\times 8$ mixture of categoricals). The VAE is optimized using the ELBO loss (Eq. 10). The Explorer is implemented as an SAC (Fig. 5(b)) and optimizes an intrinsic exploratory reward.
$\begin{split}\mathcal{L}(\text{Enc}_{U},\text{Dec}_{U})=\left\|s_{t}-\text{Dec}_{U}(z_{t})\right\|^{2}+\beta\text{KL}[\text{Enc}_{U}(z_{t}|s_{t})\parallel p_{U}(z)]\quad\text{where}\quad z_{t}\sim\text{Enc}_{U}(z_{t}|s_{t})\end{split}$ (10)
2.4.1 Exploratory Reward
Since the planning policy $\pi_{p}$ uses the GCSR decoder for subgoal prediction, it is limited to the subgoals learned by the GCSR module. The GCSR training data must contain all possible state triplets in the environment for the planner to function well. To do this, we formulate our exploratory objective to encourage the Explorer to enact trajectories not modeled well by the GCSR module. If Explorer traverses a trajectory that contains a state triplet $(s_{t},s_{t+q/2},s_{t+q})$, the modeling error is measured as the mean squared error between sub-goal $s_{t+q/2}$ and its prediction using the GCSR decoder. We compute the exploratory rewards for state $s_{t}$ based on the previous states as:
$R^{E}_{t}=\sum_{q\in Q}\left\|s_{t-q/2}-\text{Dec}_{G}(s_{t-q},s_{t},z)\right\|^{2}\quad\text{where}\quad z\sim\text{Enc}_{G}(s_{t-q},s_{t-q/2},s_{t})$ (11)
2.4.2 Memory Augmented Explorer
Since the exploratory rewards for the current step depend on the past states, the Explorer needs to know them to guide the agent accurately along rewarding trajectories. Fig. 5(c) shows the state dependencies for the exploratory rewards at different temporal resolutions $q$. Each temporal resolution $q$ requires $(s_{t-q},s_{t-q/2})$ for computing reward. For our case, the past states required are: $[(s_{t-K},s_{t-2K}),(s_{t-2K},s_{t-4K}),...]$ for each $q\in\{2K,4K,...\}$ respectively. Removing the duplicates reduces the set of states required to: $\{s_{t},s_{t-K},s_{t-2K},...\}$.
To address this, we provide a memory of the past states as an additional input to the exploratory manager SAC ($\pi_{E}(s_{t},\text{mem}_{t}),v_{E}(s_{t},\text{mem}_{t})$) (Fig. 5(b)). We implement this by maintaining a memory buffer that remembers every $K$-th visited state. Then, we extract the required states as memory for the Explorer. For a trajectory rollout of length $T$, the required size of the memory buffer is $L_{\text{mem}}=T/K$, and the size of the memory input is $\log_{2}L_{\text{mem}}$.
Practical Consideration: It can be practically infeasible to maintain a large memory buffer. However, our memory formulation is highly flexible and allows us to ignore exploratory rewards that require states far in the past. For all our experiments, we use $T=64$ length rollouts with $L_{\text{mem}}=8$ and memory input size: $3$. The trajectory length also limits the temporal resolutions $q$ for which the exploratory rewards can be computed. While not entirely optimal, this is sufficient for a significant performance improvement (Fig. 8(c)).
((a)) Unconditional VAE
((b)) Explorer subgoal prediction
((c)) Explorer input states
Figure 5: (a) Unconditional VAE that learns to predict states not conditioned on other states. (b) The explorer using the memory and the VAE decoder to predict subgoals for the worker. (c) Illustration showing the states required as inputs for the Explorer for an example trajectory (shown at the top) being played out by an agent. It is a coarse trajectory that shows every $K$-th frame. The agent is at state $s_{t}$ and will receive rewards when it moves into the placeholder future state $s_{t+1}$ (dashed border). The rewards at $s_{t+1}$ will be computed using the GCSR for different temporal resolutions $q\in Q$ indicated on the right. The colored arrows indicate the state triplet required to compute the exploratory reward at $s_{t+1}$. Combining these state dependencies and removing redundancies yields the input requirements indicated below the dashed line. The inputs consist of the current state and the memory.
2.4.3 Policy Optimization
The Explorer is optimized as an SAC using the policy gradients from the REINFORCE objective. RSSM is used to imagine trajectories $\kappa$, followed by extracting every $K$-th frame of the rollout as abstract trajectories. For clarity, let the time steps of the abstract trajectory be indexed by $k$, where $k=t/K$. Then, the lambda returns $G^{\lambda}_{k}$ are computed for the abstract trajectories using a discount factor $\gamma_{L}$ (Eq. 12). Finally, the lambda returns are used to formulate the explorer actor and critic loss (Eq. 13,14).
$G^{\lambda}_{k}=R^{E}_{k+1}+\gamma_{L}((1-\lambda)v_{E}(s_{k+1})+\lambda G^{\lambda}_{k+1})$ (12)
$\mathcal{L}(\pi_{E})=-\mathbb{E}_{\kappa\sim\pi_{E}}\sum_{k=0}^{T/K-1}(G^{\lambda}_{k}-v_{E}(s_{k},\text{mem}_{k}))\log\pi_{E}(z_{k}|s_{k},\text{mem}_{k})+\eta\mathbb{H}[\pi_{E}(z_{k}|s_{k},\text{mem}_{k})]$ (13)
$\mathcal{L}(v_{E})=\mathbb{E}_{\kappa\sim\pi_{E}}\sum_{k=0}^{T/K-1}(v_{E}(s_{k},\text{mem}_{k})-G^{\lambda}_{k})^{2}$ (14)
3 Evaluation & Results
Task Details
Figure 6: Full maze with a sample run using our agent.
We extensively test our agent in the 25-room environment, a 2D maze task where the agent has to navigate through connected rooms to reach the target position. Benchmarks from previous methods show average episode duration $>150$ steps, indicating a long-horizon task. Observations are provided as initial and goal states ($64\times 64$ images) and a reward $0<R\leq 1$ upon reaching the goal position. Each episode lasts $400$ steps before terminating with a 00 reward. We use the same evaluation parameters as the previous benchmarks and average across $100$ runs [27]. We also test our method in momentum-based environments like RoboYoga [20], which can be challenging due to the lack of momentum information in a single image (goal).
Agent Hyperparameters
We use a common hyperparameter setup for all tasks. The goal refresh rate is set to $K=8$, and the modeled temporal resolutions as $Q=\{2K,4K,8K,16K,32K\}$. The depth of the unrolled tree during training is $D=5$ and during inference is $D_{I}=8$ unless specified otherwise. For the first $3$M steps, the explorer is used as the manager, then it shifts to the planning policy. The agent is trained every $16$ environmental steps. Please refer to section B for complete training details.
3.1 Results
Figure 7 shows the sample solutions generated by our agent during training and inference. Our agent can navigate the maze to reach far goals successfully and is interpretable.
((a))
((b))
Figure 7: Samples plan by our agent during training and inference. (a) Samples of full plans as sequences of subgoals from the start states (left) to the goal states (right). The subgoals are extracted from the goal states of all terminal nodes. The blanks are when nodes terminate before maximum depth. (b) Sample subgoals generated during inference. The first and last images indicate the initial and goal states. Other images represent subgoals that break the path from the initial to the image on its right [Same order as 2(b)]. Blanks are when a reachable subgoal is found before the max depth.
We compare the performance in terms of the average success rate in reaching the goal state and the average path length against previous methods. Goal-Conditioned Behavioral Cloning (GC BC, [24]) that learns goal-reaching behavior from example goal-reaching behavior. Visual foresight (VF, [9]) that optimizes rollouts from a forward prediction model via the cross-entropic method (CEM, [29, 23]). Hierarchical planning using Goal-Conditioned Predictors (GCP,[27]) optimized using CEM to minimize the predicted distance cost. Goal-Conditioned Director (GC-Director). And LEXA, a SOTA sequential planning method that also uses an explorer and a planner but optimizes continuous rewards cosine and temporal. GC BC, VF, and GCP performances are taken from [27], which uses the same evaluation strategy. Table 1 shows that our model outperforms the previous work in terms of success rate and average episode lengths. Our method and LEXA (Temp) yield the shortest episode lengths. However, note that while LEXA explicitly optimizes for path lengths, our method uses an implicit objective.
Agent Success rate Average episode length Sample complexity Time complexity
GC BC $7\%$ $402.48$ $1$ $1$
GC-Director $9\%$ $378.89\pm 87.67$ $1$ $1$
LEXA (Cos) $20\%$ $321.04\pm 153.29$ $1$ $1$
VF $26\%$ $362.82$ $MN$ $N$
GCP $82\%$ $158.06$ $(2N+1)M$ $\log N$
LEXA (Temp) $90\%$ $70.34\pm 111.14$ $1$ $1$
DHP (Ours) $100\%$ $73.84\pm 46.54$ $\log N$ $\log N$
Table 1: Average Performance of different approaches on the 25-room navigation task over $100$ evaluation runs. $N$ is the number of plan steps, and $M$ is the number of samples generated per plan.
Fig. 8(a) shows the score and episode length plots for some methods. Here, we plot an extra experiment, Director (Fixed Goal), where the goal remains fixed and the agent only inputs the current state image. It can be seen that the agent shows signs of learning, the score falls but rises again around $8$M steps to $\sim 70\%$. Comparing this to GC-Director (which completely fails) shows that the issue is not navigation or agent size, but the complexity of a goal-conditioned long-horizon task that requires planning.
((a)) Baselines
((b)) Bootstapping
((c)) Exploration
((d)) Other rewards schemes
Figure 8: Experimental results (mean and stddev across $3$ seeds) shown with episodic reward (top row) and average episode length (bottom row) against the environmental steps. The sharp performance rise for our methods indicates the switch from exploration to planning. (a) Baselines: Comparing our method DHP with state-of-the-art methods. (b) Bootstrapping: Compares shallow training depth models $(D=1,3)$ with the default training depth. (c) Explore data Comparison of planning policy performance trained on different data. (d) Comparison of Other reward schemes.
3.2 Ablations
Can the planning method generalize to higher depths?
We train two DHP agents with maximum tree depth during planning to $3$ and $1$. The $1$ depth agent is only allowed to break the given task once for learning. Fig. 8(b) shows the comparison of the agents, and it can be seen that all agents perform similarly.
Does the complex exploration strategy help?
We compare the performance of the planning policy trained against three variants: using expert data (collected using a suboptimal policy in [27]), vanilla exploration objective using the reconstruction errors from the unconditional VAE (Sec. 2.4), and using an explorer without the memory. Fig. 8(c) shows the comparison, and it can be seen that the default agent performs noticeably better.
Can the method perform in other environments?
Figure 9: Robo Yoga Environments
We test our agent in the Deepmind Control Suite [34] based RoboYoga [20] environment. The task requires walker and quadruped agents to reach a goal body orientation specified by an image. The environment is challenging because it rewards the agent to maintain the goal position, while our agent only plans to reach the goal state. Also, the goal states do not encode the momentum information. Fig. 9 shows the performance, of our agent at the tasks. Our agent can solve the walker task, but struggles at quadruped. We observed that our walker agent maintains the overall pose but constantly sways about the goal position. Fig. 13 shows an episode where the agent headstands constantly.
How does the method perform with other reward and return schemes?
We compare the default agent against a few variants: DHP (Neg Rew) That rewards $-1$ at non-terminal nodes and 00 at terminal (like an existence penalty), DHP (GAE) that estimates advantages as $\min$ of the GAE advantages for each child node, and DHP (Dist Sum) rewards negative estimated distances (between $n_{i,0}$ and $n_{i,1}$) at each node. Fig. 8(d) shows that the Neg Rew agent performs as well as the default method, while the others don’t perform as well.
4 Related Work
Hierarchical RL Agents: Hierarchical reinforcement learning (HRL) is a set of techniques that abstracts actions [4, 35, 3, 32, 26, 21]. Foundational works like the options framework [32] and MAXQ decomposition [8] introduced temporal abstraction, allowing agents to reason at multiple time scales. Modern approaches learn hierarchical policies through mutual information (Causal InfoGAN [17], DADS [30]), latent space planning (Director [14]), or trajectory encoding (OPAL [1]). These show that hierarchical decomposition aids efficient exploration and credit assignment for planning.
Planning Algorithms: Planning methods aim to solve long-horizon tasks efficiently by exploring future states and selecting optimal actions [18, 6]. Monte Carlo Tree Search (MCTS) [5] expands a tree of possible future states by sampling actions and simulating outcomes. While effective in discrete action spaces, MCTS struggles with scalability in high-dimensional or continuous environments. Visual Foresight methods [9, 11, 13] learned visual dynamics models to simulate future states, enabling planning in pixel-based environments. However, they require accurate world models and can be computationally expensive. Some use explicit graph search over the replay buffer data [10]. Model Predictive Control (MPC) [22, 23] is an online planner that samples future trajectories and optimizes actions over a finite horizon. These methods rely on sampling the future state and thus do not scale well with the horizon length. LEXA [16] is a policy-based linear planner that also uses an explorer for data collection.
To address the challenges of long-horizon tasks, some planning algorithms decompose complex problems into manageable subtasks by predicting intermediate subgoals [25, 15, 27]. MAXQ [8] decomposes the value function of the target Markov Decision Process (MDP) into an additive combination of smaller MDPs. Sub-Goal Trees [15] and CO-PILOT [2] learn a subgoal prediction policy optimized to minimize the total predicted distance measures of the decomposed subtasks. GCP [27] and kSubS [7] use specialized subgoal predicting modules to search for plans. DHRL [19] uses efficient graph search by decoupling the time horizons of high-level and low-level policies.
While these methods have demonstrated success, they rely heavily on distance-based metrics, which are challenging to learn and sensitive to policy quality [10, 2]. In contrast, our method uses discrete reachability-based rewards that are easier to accurately estimate and provide clearer learning signals.
5 Discussion & Future Work
DHP architecture allows us to train a future-conditioned (goal) planning policy $\pi_{P}$ and a past-conditioned (memory) explorer policy $\pi_{E}$. The resulting model performs expertly on the standard 25-room long-horizon task, where the current SOTA approaches fail (Fig. 8(a)). The ablations show that the method generalizes beyond the training depths (Fig. 8(b)), the exploratory rewards significantly impact performance which are further enhanced using memory (Fig. 8(c)), and possible alternate reward schemes (Fig. 8(d)). Our method relies on imagination for training, which allows plan evaluation without the need to access the internal environmental state, but it may have inaccuracies (Fig. 9). However, we observed better results with discrete rewards than with the distance-based approach for our model (Fig. 8(d)). The architecture is flexible, and components can be used in isolation, e.g., the planning policy optimization can be combined with custom reachability measures. For future work, the agent can benefit from better and generic goal-state estimation mechanisms, preferably multimodal. The agent can also learn to generate goals for itself via a curriculum for targeted exploration, which can benefit exploration in complex environments with multiple bottleneck states that require a long-term commitment to a goal for efficient exploration. Given this, we show that the agent can be helpful at solving long-horizon tasks and believe that the ideas presented in the paper can be beneficial to the community. The code for the agent is available at: <URL redacted>.
Appendix A Theoretical Proofs and Derivations
A.1 Policy Gradients for Trees
Given a goal-directed task as a pair of initial and final states $(s_{t},s_{g})$, a subgoal generation method predicts an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator further breaks the task, leading to a tree of subtasks $\tau$ where each node $n_{i}$ is a task. Let the preorder traversal of the subtask tree $\tau$ of depth $D$ be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. The root node $n_{0}$ is the given task and the other nodes are the recursively generated subtasks. Ideally, the leaf nodes should indicate the simplest reduction of the subtask that can be executed sequentially to complete the original task. The tree can be viewed as a trajectory where each node $n_{i}$ is a state, and taking action $\pi_{P}(z_{i}|n_{i})$ simultaneously places the agent in two states given by the child nodes $(n_{2i+1},n_{2i+2})$. Thus, the policy function can be written as $\pi_{P}(z_{i}|n_{i})$, the transition probabilities (can be deterministic also) as $p_{T}(n_{2i+1},n_{2i+2}|z_{i},n_{i})$, and the probability of the tree trajectory under the policy $\tau_{\pi_{P}}$ can be represented as:
$p_{\pi_{P}}(\tau)$ $=p(n_{0})*[\pi_{P}(z_{0}|n_{0})*p_{T}(n_{1},n_{2}|z_{0},n_{0})]*$
$\quad[\pi_{P}(z_{1}|n_{1})*p_{T}(n_{3},n_{4}|z_{1},n_{1})]*$
$\quad[\pi_{P}(z_{2}|n_{2})*p_{T}(n_{5},n_{6}|z_{2},n_{2})]*...$
$\quad[\pi_{P}(z_{2^{D}-2}|n_{2^{D}-2})*p_{T}(n_{2^{D+1}-3},n_{2^{D+1}-2}|z_{2^{D}-2}|n_{2^{D}-2})]$
$p_{\pi_{P}}(\tau)$ $=p(n_{0})\prod_{i=0}^{2^{D}-2}\pi_{P}(z_{i}|n_{i})\prod_{i=0}^{2^{D}-2}p_{T}(n_{2i+1},n_{2i+2}|z_{i},n_{i})$
Theorem A.1 (Policy Gradients).
Given a tree trajectory $\tau$ specified as a list of nodes $n_{i}$, generated using a policy $\pi_{P}$. The policy gradients can be written as:
$\nabla_{\pi_{P}}J({\pi_{P}})=\mathbb{E}_{\tau}\sum_{i=0}^{2^{D}-2}A^{i}(\tau)\nabla_{\pi_{P}}\log\pi_{\pi_{P}}(z_{i}|n_{i})$
Proof.
The log-probabilities of the tree trajectory and their gradients can be written as:
$\log p_{\pi_{P}}(\tau)=\log p(n_{0})+\sum_{i=0}^{2^{D}-2}\log\pi_{P}(z_{i}|n_{i})+\sum_{i=0}^{2^{D}-2}\log p_{T}(n_{2i+1},n_{2i+2}|z_{i},n_{i})$ (15)
$\nabla_{\pi_{P}}\log p_{\pi_{P}}(\tau)=0+\nabla_{\pi_{P}}\sum_{i=0}^{2^{D}-2}\log\pi_{P}(z_{i}|n_{i})+0=\sum_{i=0}^{2^{D}-2}\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})$ (16)
The objective of policy gradient methods is measured as the expectation of advantage or some scoring function $A(\tau)$:
$J(\pi_{P})=\mathbb{E}_{\tau}A(\tau)=\sum_{\tau}A(\tau)\cdot p_{\pi_{P}}(\tau)$ (17)
Then the gradients of the objective function $\nabla_{\pi_{P}}J(\pi_{P})$ wrt the policy $\pi_{P}$ can be derived as:
$\nabla_{\pi_{P}}J(\pi_{P})$ $=\nabla_{\pi_{P}}\sum_{\tau}A(\tau)\cdot p_{\pi_{P}}(\tau)$
$=\sum_{\tau}A(\tau)\cdot\nabla_{\pi_{P}}p_{\pi_{P}}(\tau)$
$=\sum_{\tau}A(\tau)\cdot p_{\pi_{P}}(\tau)\frac{\nabla_{\pi_{P}}p_{\pi_{P}}(\tau)}{p_{\pi_{P}}(\tau)}$
$=\sum_{\tau}A(\tau)\cdot p_{\pi_{P}}(\tau)\nabla_{\pi_{P}}\log p_{\pi_{P}}(\tau)$
$=\mathbb{E}_{\tau}A(\tau)\cdot\nabla_{\pi_{P}}\log p_{\pi_{P}}(\tau)$
[…]
Theorem A.2 (Baselines).
If $A(\tau)$ is any function independent of policy actions $z_{i}$, say $b(n_{i})$, then its net contribution to the policy gradient is 00.
$\mathbb{E}_{\tau}\sum_{i=0}^{2^{D}-2}b(n_{i})\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})=0$
Proof.
If $A(\tau)$ is any fixed function that does not depend on the actions $\pi_{P}(z_{i}|n_{i})$ and only on the state, say $b(n_{i})$. Then $b(n_{i})$ will be independent of the trajectory $\tau$, and it can be sampled from the steady state distribution under policy $\rho_{\pi_{P}}$ for any state $n_{i}$ without knowing $\tau$. In that case,
$\mathbb{E}_{\tau}\sum_{i=0}^{2^{D}-2}b(n_{i})\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})$ $=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{\tau}[b(n_{i})\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})]$
$=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{n_{i}\sim\rho_{\pi_{P}}}\mathbb{E}_{a\sim\pi_{P}}[b(n_{i})\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})]$
$=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{n_{i}\sim\rho_{\pi_{P}}}[b(n_{i})\mathbb{E}_{a\sim\pi_{P}}\nabla_{\pi_{P}}\log\pi_{P}(z_{i}|n_{i})]$
$=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{n_{i}\sim\rho_{\pi_{P}}}[b(n_{i})\sum_{a}\pi_{P}(z_{i}|n_{i})\frac{\nabla_{\pi_{P}}\pi_{P}(z_{i}|n_{i})}{\pi_{P}(z_{i}|n_{i})}]$
$=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{n_{i}\sim\rho_{\pi_{P}}}b(n_{i})[\nabla_{\pi_{P}}\sum_{a}\pi_{P}(z_{i}|n_{i})]$
$=\sum_{i=0}^{2^{D}-2}\mathbb{E}_{n_{i}\sim\rho_{\pi_{P}}}b(n_{i})[\nabla_{\pi_{P}}1]\quad(\text{Sum of probabilities is 1})$
$=0$
A.2 Policy Evaluation for Trees
We present the return and advantage estimation for trees as an extension of current return estimation methods for linear trajectories. As the return estimation for a state $s_{t}$ in linear trajectories depends upon the next state $s_{t+1}$, our tree return estimation method uses child nodes $(n_{2i+1},n_{2i+2})$ to compute the return for a node $n_{i}$. We extend the previous methods like lambda returns and Gen realized Advantage estimation (GAE) for trees.
The objective of our method is to reach nodes that are directly reachable. Such nodes are marked as terminal, and the agent receives a reward. For generalization, let’s say that when the agent takes an action $z_{i}$ at node $n_{i}$, it receives a pair rewards $(R_{2i+1}(n_{i},z_{i}),R_{2i+2}(n_{i},z_{i}))$ corresponding to the child nodes. Formally, the rewards $R(\tau)$ is an array of length equal to the length of the tree trajectory with $R_{0}=0$. Then, the agent’s task is to maximize the sum of rewards received in the tree trajectory $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}R_{i}$. To consider future rewards, the returns for a trajectory can be computed as the sum of rewards discounted by their distance from the root node (depth), $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}\gamma^{\lfloor\log_{2}(i+1)\rfloor-1}R_{i}$. Thus, the returns for each node can be written as the sum of rewards obtained and the discount-weighted returns thereafter:
$G_{i}=(R_{2i+1}+\gamma G_{2i+1})+(R_{2i+2}+\gamma G_{2i+2})$ (18)
Although this works theoretically, a flaw causes the agent to collapse to a degenerate local optimum. This can happen if the agent can generate a subgoal very similar to the initial or goal state ($\|s_{t},s_{\text{sub}}\|<\epsilon$ or $\|s_{g},s_{\text{sub}}\|<\epsilon$). A usual theme of reward systems for subgoal trees will be to have a high reward when the agent predicts a reachable or temporally close enough subgoal. Thus, if the agent predicts a degenerate subgoal, it receives a reward for one child node and the initial problem carries forward to the other node.
Therefore, we propose an alternative objective that optimizes for the above objective under the condition that both child subtasks $(n_{2i+1},n_{2i+2})$ get solved. Instead of estimating the return as the sum of the returns from the child nodes, we can estimate it as the minimum of the child node returns.
$G_{i}=\min(R_{2i+1}+\gamma G_{2i+1},R_{2i+2}+\gamma G_{2i+2})$ (19)
This formulation causes the agent to optimize the weaker child node first and receive discounted rewards if all subtasks are solved (or have high returns). It can also be noticed that the tree return for a node is essentially the discounted return along the linear trajectory that traces the path with the least return starting at that node. Next, we analyze different return methods in the tree setting and try to prove their convergence.
A.2.1 Lambda Returns
TD($\lambda$) returns for linear trajectories are computed as:
$G_{t}^{\lambda}=R_{t+1}+\gamma((1-\lambda)V(s_{t+1})+\lambda G_{t+1}^{\lambda})$ (20)
We propose, the lambda returns for tree trajectories can be computed as:
$G_{i}^{\lambda}=\min(R_{2i+1}+\gamma((1-\lambda)V(n_{2i+1})+\lambda G_{2i+1}^{\lambda}),R_{2i+2}+\gamma((1-\lambda)V(s_{2i+2})+\lambda G_{2i+2}^{\lambda}))$ (21)
Which essentially translates to the minimum of lambda returns using either of the child nodes as the next state. For theoretical generalization, note that the $\min$ operator in the return estimate is over the next states the agent is placed in. Thus, in the case of linear trajectories where there is only one next state, the $\min$ operator vanishes and the equation conveniently reduces to the standard return formulation for linear trajectories.
Next, we check if there exists a fixed point that the value function approaches. The return operators can be written as:
$\mathcal{T}^{\lambda}V(n_{i})=\mathbb{E}_{\pi}[\min($ $R_{2i+1}+\gamma((1-\lambda)V(n_{2i+1})+\lambda G_{2i+1}),$
$R_{2i+2}+\gamma((1-\lambda)V(n_{2i+2})+\lambda G_{2i+2}))]$
$\mathcal{T}^{0}V(n_{i})=\mathbb{E}_{\pi}[\min($ $R_{2i+1}+\gamma V(n_{2i+1}),R_{2i+2}+\gamma V(n_{2i+2}))]$
$\mathcal{T}^{1}V(n_{i})=\mathbb{E}_{\pi}[\min($ $R_{2i+1}+\gamma G_{2i+2},R_{2i+2}+\gamma G_{2i+2})]$
Lemma A.3 (Non-expansive Property of the Minimum Operator).
For any real numbers $a,b,c,d$, the following inequality holds:
$\left|\min(a,b)-\min(c,d)\right|\leq\max\left(|a-c|,|b-d|\right).$
Proof.
To prove the statement, we consider the minimum operator for all possible cases of $a$, $b$, $c$, and $d$. Let $\min(a,b)$ and $\min(c,d)$ be the minimum values of their respective pairs.
Case 1: $a\leq b$ and $c\leq d$
In this case, $\min(a,b)=a$ and $\min(c,d)=c$. The difference becomes:
$\left|\min(a,b)-\min(c,d)\right|=|a-c|.$
Since $\max\left(|a-c|,|b-d|\right)\geq|a-c|$, the inequality holds.
Case 2: $a\leq b$ and $c>d$
Here, $\min(a,b)=a$ and $\min(c,d)=d$. The difference becomes:
$\left|\min(a,b)-\min(c,d)\right|=|a-d|.$
Since $b\geq a$, $|a-d|\leq|b-d|\leq\max(|a-c|,|b-d|)$, and the inequality holds.
Case 3: $a>b$ and $c\leq d$
Here, $\min(a,b)=b$ and $\min(c,d)=c$. The difference becomes:
$\left|\min(a,b)-\min(c,d)\right|=|b-c|.$
Since $a\geq b$, $|b-c|\leq|a-c|\leq\max(|a-c|,|b-d|)$, and the inequality holds.
Case 4: $a>b$ and $c>d$
Symmetric to Case 1.
Conclusion:
In all cases, the inequality
$\left|\min(a,b)-\min(c,d)\right|\leq\max\left(|a-c|,|b-d|\right)$
is satisfied. Therefore, the minimum operator is non-expansive. ∎
Theorem A.4 (Contraction property of the return operators).
The Bellman operators $\mathcal{T}$ corresponding to the returns are a $\gamma$-contraction mapping wrt. to $\|\cdot\|_{\infty}$
$\|\mathcal{T}V_{1}-\mathcal{T}V_{2}\|_{\infty}\leq\gamma\|V_{1}-V_{2}\|_{\infty}$
Proof.
We start with the simpler case, $\mathcal{T}^{0}$. Let $V_{1},V_{2}$ be two arbitrary value functions. Then the max norm of any two points in the value function post update is:
$\|\mathcal{T}^{0}V_{1}-\mathcal{T}^{0}V_{2}\|_{\infty}=$ $\|\mathbb{E}_{\pi}[\min(R_{2i+1}+\gamma V_{1}(n_{2i+1}),R_{2i+2}+\gamma V_{1}(n_{2i+2}))]-$
$\mathbb{E}_{\pi}[\min(R_{2i+1}+\gamma V_{2}(n_{2i+1}),R_{2i+2}+\gamma V_{2}(n_{2i+2}))]\|_{\infty}$
$=$ $\|\mathbb{E}_{\pi}[\min(R_{2i+1}+\gamma V_{1}(n_{2i+1}),R_{2i+2}+\gamma V_{1}(n_{2i+2}))-$
$\min(R_{2i+1}+\gamma V_{2}(n_{2i+1}),R_{2i+2}+\gamma V_{2}(n_{2i+2}))]\|_{\infty}$
$\leq$ $\|\min(R_{2i+1}+\gamma V_{1}(n_{2i+1}),R_{2i+2}+\gamma V_{1}(n_{2i+2}))-$
$\min(R_{2i+1}+\gamma V_{2}(n_{2i+1}),R_{2i+2}+\gamma V_{2}(n_{2i+2}))\|_{\infty}$
$\leq$ $\max(\|\gamma V_{1}(n_{2i+1})-\gamma V_{2}(n_{2i+1})\|_{\infty},\|\gamma V_{1}(n_{2i+2})-\gamma V_{2}(n_{2i+2})\|_{\infty})$
$\leq$ $\gamma\max(\|V_{1}(n_{2i+1})-V_{2}(n_{2i+1})\|_{\infty},\|V_{1}(n_{2i+2})-V_{2}(n_{2i+2})\|_{\infty})$
$\leq$ $\gamma\max(\|V_{1}(n_{j})-V_{2}(n_{j})\|_{\infty},\|V_{1}(n_{k})-V_{2}(n_{k})\|_{\infty})$
$\leq$ $\gamma\|V_{1}-V_{2}\|_{\infty}\quad\text{(merging $\max$ with $\|\cdot\|_{\infty}$)}$
A similar argument can be shown for $\|\mathcal{T}^{1}V_{1}-\mathcal{T}^{1}V_{2}\|_{\infty}$ and $\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}$. Using the non-expansive property (Th. A.3) and absorbing the $\max$ operator with $\|\cdot\|_{\infty}$ leading to the standard form for linear trajectories.
$\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}\leq$ $\|\gamma((1-\lambda)(V_{1}-V_{2})+\lambda(\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}))\|_{\infty}$
[…]
$(1-\gamma\lambda)\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}\leq$ $\gamma(1-\lambda)\|V_{1}-V_{2}\|_{\infty}$
$\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}\leq$ $\frac{\gamma(1-\lambda)}{1-\gamma\lambda}\|V_{1}-V_{2}\|_{\infty}$
For contraction, $\frac{\gamma(1-\lambda)}{1-\gamma\lambda}<1$ must be true.
$\frac{\gamma(1-\lambda)}{1-\gamma\lambda}<$ 11 11
$\gamma(1-\lambda)<$ ${1-\gamma\lambda}$
$\gamma-\gamma\lambda<$ $1-\gamma\lambda$
$\gamma<$ 11 11
Which is always true.
Since $\mathcal{T}^{1}$ is a special case of $\mathcal{T}^{\lambda}$, it is also a contraction. ∎
A.2.2 Bootstrapping with $D$-depth Returns
When the subtask tree branches end as terminal (or are masked as reachable), the agent receives a reward $=1$, which provides a learning signal using the discounted returns. However, when the branches do not end as terminal nodes, it does not provide a learning signal for the nodes above it, as the return is formulated as $\min$ of the returns from child nodes. In this case, we can replace the returns of the non-terminal leaf nodes with their value estimates. Therefore, in case the value estimate from the end node is high, indicating the agent knows how to solve the task from there onwards, it still provides a learning signal. The $n$-step return for a linear trajectory is written as:
$G^{(n)}_{t}=R_{t+1}+\gamma G^{(n-1)}_{t+1}$
with the base case as:
$G_{t}^{(1)}=R_{t+1}+\gamma V(s_{t+1})$
We write the $n$-step returns for the tree trajectory as:
$G_{i}^{(d)}=\min(R_{2i+1}+\gamma G^{(d-1)}_{2i+1},R_{2i+2}+\gamma G^{(d-1)}_{2i+2})$
with the base case as:
$G_{i}^{(1)}=\min(R_{t+1}+\gamma V(n_{2i+1}),R_{t+2}+\gamma V(n_{2i+2}))$
Value estimates help bootstrap at the maximum depth of the unrolled subtask tree $D$ and allow the policy to learn from incomplete plans.
A.2.3 Properties of Tree Return Estimates
In section A.2 it can be seen how the tree return formulation for a node essentially reduces to the linear trajectory returns along the path of minimum return in the subtree under it. When the value function has reached the stationary point. For a subtask tree, if all branches end as terminal nodes, the return will be $\gamma^{D^{\prime}}1$, where $D^{\prime}$ is the depth of the deepest non-terminal node. Otherwise, it would be $\gamma^{D}V^{\prime}$ where $V^{\prime}$ is the non-terminal truncated node with the minimum return. Thus, it can be seen how higher depth penalizes the returns received at the root node with the discount factor $\gamma$. This property is true with the linear trajectories where the policy converges to the shortest paths to rewards to counter discounting [31, 28]. Thus, our goal-conditioned policy similarly converges to plans trees with minimum maximum depth: $\min_{\pi_{P}}(\max d_{i})$, where $d_{i}$ is the depth of node $n_{i}$.
This property also implies that the returns for a balanced tree will be higher than those for an unbalanced tree. The same sequence of leaf nodes can be created using different subtask trees. When the policy does not divide the task into roughly equally tough sub-tasks it results in an unbalanced tree. Since the tree is constrained to yield the same sequence of leaf nodes, its maximum depth $D_{U}$ will be higher than or equal to a balanced tree $D_{U}\geq D_{B}$. Thus, at optimality, the policy should subdivide the task in roughly equal chunks. However, it should be noted that two subtask trees with different numbers of leaf nodes can have the same maximum depth.
Appendix B Architecture & Training Details
B.1 Worker
The worker is trained using $K$-step imagined rollouts $(\kappa\sim\pi_{W})$. Given the imagined trajectory $\kappa$, the rewards for the worker $R^{W}_{t}$ are computed as the cosine_max similarity measure between the trajectory states $s_{t}$ and the prescribed worker goal $s_{\text{wg}}$. First, discounted returns $G^{\lambda}_{t}$ are computed as $n$-step lambda returns (Eq. 23). Then the Actor policy is trained using the REINFORCE objective (Eq. 24) and the Critic is trained to predict the discounted returns (Eq. 25).
$R^{W}_{t}=\texttt{cosine\_max}(s_{t},s_{\text{wg}})$ (22)
$G^{\lambda}_{t}=R^{W}_{t+1}+\gamma_{L}((1-\lambda)v(s_{t+1})+\lambda G^{\lambda}_{t+1})$ (23)
$\mathcal{L}(\pi_{W})=-\mathbb{E}_{\kappa\sim\pi_{W}}\sum_{t=0}^{K-1}\left[(G^{\lambda}_{t}-v_{W}(s_{t}))\log\pi_{W}(z|s_{t})+\eta\text{H}[\pi_{W}(z|s_{t})]\right]$ (24)
$\mathcal{L}(v_{W})=\mathbb{E}_{\kappa\sim\pi_{W}}\sum_{t=0}^{K-1}(v_{W}(s_{t})-G^{\lambda}_{t})^{2}$ (25)
B.2 Goal State Representations
Since the RSSM integrates a state representation using a sequence of observations, it does not work well for single observations. To generate goal state representations using single observations we train an MLP separately that tries to approximate the RSSM outputs ($s_{t}$) from the single observations ($o_{t}$). We call these representations static state representations. Moreover, since GCSR modules require state representations at large temporal distances, it can be practically infeasible to generate them using RSSM. Thus, we use static state representations to generate training data for the GCSR module as well. The MLP is a dense network with a $\tanh$ activation at the final output layer. It is trained to predict the RSSM output (computed using a sequence of images) using single image observations. To avoid saturating gradients, we use an MSE loss on the preactivation layer using labels transformed as $l_{\text{new}}=\text{atanh}(\text{clip}(l,\delta-1,1-\delta))$. The clipping helps avoid computational overflows, we use $\delta=10^{-4}$.
B.3 Implementation Details
We implement two functions: policy and train, using the hyperparameters shown in Table 2. The agent is implemented in Python/Tensorflow with XLA JIT compilation. Using XLA optimizations, the total training wall time is $2-3$ days on a consumer GPU (NVIDIA 4090 RTX 24gb). During inference, the compiled policy function runs in $4.6$ms on average, enabling real-time replanning at $217.39$ times per second (assuming minimal environmental overhead).
Name Symbol Value
Train batch size $B$ $16$
Replay trajectory length $L$ $64$
Replay coarse trajectory length $L_{c}$ $48$
Worker abstraction length $K$ $8$
Explorer Imagination Horizon $H_{E}$ $64$
Return Lambda $\lambda$ $0.95$
Return Discount (tree) $\gamma$ $0.95$
Return Discount (worker & explorer) $\gamma_{L}$ $0.99$
State similarity threshold $\Delta_{R}$ $0.7$
Plan temporal resolutions $Q$$\{16,32,64,128,256\}$
Maximum Tree depth during training $D$ $5$
Maximum Tree depth during inference $D_{\text{Inf}}$ $8$
Target entropy $\eta$ 0.5
KL loss weight $\beta$ 1.0
GCSR latent size - $4\times 4$
RSSM deter size - $1024$
RSSM stoch size - $32\times 32$
Optimizer - Adam
Learning rate (all) - $10^{-4}$
Adam Epsilon - $10^{-6}$
Optimizer gradient clipping - 1.0
Weight decay (all) - $10^{-2}$
Activations - LayerNorm + ELU
MLP sizes - $4\times 512$
Train every - 16
Prallel Envs - 1
Table 2: Agent Hyperparameters
B.3.1 Policy Function
At each step, the policy function is triggered with the environmental observation $o_{t}$. The RSSM module processes the observation $o_{t}$ and the previous state $s_{t-1}$ to yield a state representation $s_{t}$. During exploration, the manager $\pi_{E}$ uses the $s_{t}$ to generate a worker goal using the unconditional VAE. During task policy, the planning manager $\pi_{P}$ generates subgoals in the context of a long-term goal $s_{g}$, and the first directly reachable subgoal is used as the worker’s goal. Finally, the worker generates a low-level environmental action using the current state and the worker goal $(s_{t},s_{\text{wg}})$. The algorithm is illustrated in (Alg. 1)
Algorithm 1 DHP Policy Function
0:  Observation $o_{t}$, Goal $o_{g}$, agent state $t$, $s_{t-1}$, $a_{t-1}$, $s_{wg}$, mem_buff, mode
0:  Action $a_{t}$, new agent state
1:  $s_{t}\leftarrow\texttt{wm}(o_{t},s_{t-1},a_{t-1})$World model update}
2:  if $t\mod K$ = 0 then
3:     if mode == $eval$ then
4:        $s_{g}\leftarrow$ static_state($o_{g}$) {Sec B.2}
5:        $d\leftarrow 0$Subgoal planning}
[…]
7:           $z\sim\pi_{P}(z|s_{t},s_{g})$
8:           $s_{g}\leftarrow\text{Dec}_{G}(z,s_{t},s_{g})$
9:           $d\leftarrow d+1$
10:        end while
11:        $s_{wg}\leftarrow s_{g}$
12:     else
[…]
14:        $z\sim\pi_{E}(z|s_{t},\text{mem}_{t})$
15:        $s_{wg}\leftarrow\text{Dec}_{U}(z)$
16:     end if
[…]
18:  end if
19:  $a_{t}\sim\pi_{W}(a_{t}|s_{t},s_{wg})$
[…]
B.3.2 Train Function
The training function is executed every $16$-th step. A batch size $B$ of trajectories $\kappa$ and coarse trajectories $\kappa_{c}$ is sampled from the exploration trajectories or the expert data. The length of extracted trajectories is $L$ and the length of coarse trajectories is $L_{c}$ spanning over $L_{c}\times K$ time steps. Then the individual modules are trained sequentially (Alg. 6):
  • RSSM module is trained using $\kappa$ via the original optimization objective [13] followed by the static state representations (Sec. B.2).
  • The GCSR module is trained using the coarse trajectory $\kappa_{c}$ (Sec. 2.2).
  • The worker policy is optimized by extracting tuples $(s_{t},s_{t+K})$ from the trajectories $\kappa$ and running the worker instantiated at $s_{t}$ with worker goal as $s_{t+K}$ (Sec. B.1).
  • The planning policy is trained using sample problems extracted as pairs of initial and final states $(s_{t},s_{g})$ at randomly mixed lengths from $\kappa_{c}$. Then the solution trees are unrolled and optimized as in Sec. 2.3.
  • Lastly, the exploratory policy is also optimized using each state in $\kappa$ as the starting state (Sec. 2.4.3).
Algorithm 2 Training the GCSR Module
0:  Experience dataset $data$
1:  Initialize $\text{triplets}\leftarrow\emptyset$
2:  for all window sizes $q\in Q$ do
[…]
4:  end for
[…]
Algorithm 3 Training the Worker Policy
0:  Experience dataset $data$
[…]
2:  $\text{traj}\leftarrow\text{imagine}(\text{init},K)$On-policy trajectory rollout}
[…]
Algorithm 4 Training the Planner Policy
0:  Experience dataset $data$
[…]
Algorithm 5 Training the Explorer Policy
0:  Experience dataset $data$
[…]
2:  $\text{traj}\leftarrow\text{imagine}(\text{init},H_{E})$On-policy trajectory rollout}
[…]
Algorithm 6 Overall Training Procedure
0:  Experience dataset $data$, current mode $mode$
[…]
Appendix C Broader Impacts
C.1 Positive Impacts
The imagination-based policy optimization mitigates hazards that can occur during learning. The efficient training can reduce the carbon footprint of the agents. The agent produces highly interpretable plans that can be verified before execution.
C.2 Negative Impacts and Mitigations
  • Inaccurate Training: Imagination can cause incorrect learning. Mitigation: Rigorous testing using manual verification of world-model reconstructions against ground truths.
  • Malicious Use: Hierarchical control could enable more autonomous adversarial agents. Mitigation: Advocate for gated release of policy checkpoints.
C.3 Limitations of Scope
Our experiments focus on simulated tasks without human interaction. Real-world impacts require further study of reward alignment and failure modes.
Appendix D Sample Exploration Trajectories
We compare sample trajectories generated by various reward schemes in this section. Fig. 10 shows the sample trajectories of an agent trained to optimize the vanilla exploratory rewards. Fig. 11 shows the sample exploration trajectories of an agent that optimizes only for the GCSR-based rewards. Fig. 12 shows sample exploration trajectories of an agent that optimizes for GCSR exploratory rewards but without memory. The GCSR rewards-based agent generates trajectories that are less likely to lead to repeated path segments or remaining stationary. Removing the memory sometimes causes inefficient trajectories where the agent might not perform as expected.
Figure 10: Sample exploration trajectories using the vanilla exploratory rewards. The agent identifies states with unclear representations and successfully navigates to them, indicating sufficient navigation capabilities. However, once at the goal, the agent stays there, leading to lower-quality data.
Figure 11: Sample exploration trajectories using the GCSR modules for the path-segments-based rewards. It can be seen that the agent continuously moves and explores longer state connectivity.
Figure 12: Sample exploration trajectories using the default strategy without the Memory. The resulting data can have some problematic trajectories.
Appendix E Robo Yoga Samples
Figure 13: Figure shows a sample trajectory (every $8$-th frame) for the walker embodiment given a headstand goal. The agent can maintain a constant headstand; however, it sways about the goal position. This is because the agent is trained to reach the goal position and not to stay there.
Figure 14: Some failure cases from the plans constructed for the yoga quadruped task where the agent hallucinates wrong midway states that get verified by the similarity check. We believe this is because the quadruped agent has a richer state space than the walker, and the states are not very distinct.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction make precise claims that align with the technical contributions and empirical validation:
    • Discrete rewards scheme performance (Fig. 8(a)).
    • Advantage estimation that generalizes beyond training depths with theoretical analysis (Fig. 8(b), Sec. A.2.3)
    • Exploration reward that encourages relevant trajectories (Fig. 8(c),11).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: While we observe that our method succeeds in challenging visual goal conditioned planning scenarios, we have observed that it struggles in certain environments, like quadruped, and we have provided this result. Thus, our method may be suitable for certain environments, but not for all, and the related results have been provided.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: All theoretical results include explicit assumptions and complete proofs (Sec. A).
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We include the full training details in the section B that includes all the important hyperparameters, architecture, and optimization objectives for reproducing the results. Although our implementation is based on Director [14], all novel components are self-contained in the algorithms provided (Algs. 1 6). Code release will include a Dockerfile for environment replication and pre-trained checkpoints.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: While we commit to releasing code post-acceptance, the current submission lacks:
    • Anonymized code/scripts in supplemental materials
    • Environment setup instructions (Dockerfile)
    • Exact reproduction commands for key experiments
    Post-release will include:
    • Complete code with environment setup.
    • Pre-trained checkpoints for all tasks.
    • Jupyter notebooks for all results and visualizations.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We try to provide all the details specific to our approach in the main paper. We include the remaining details in the supplementary (Sec. B), which include the hyperparameters and the function algorithms implemented.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: The statistical descriptions have been added to the paper. For all our plots, the results are mean and standard deviations (std) across $3$ seeds. The bar graphs are shown as mean and std over $100$ evaluation runs. However, we have a small sample size of $3$ seeds due to resource constraints. Error bars represent population standard deviation (not SEM) calculated via numpy.ndarray.std(). We plan $5$-seed runs for experiments for the camera-ready version.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. We include the exact implementation and resource details (time and compute) in the section B.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We do not include any information that can violate anonymity. We try to cite all references to previous work and sources.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss the possible impact of our work in section C. We emphasize that these are speculative, given our simulation-only experiments.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We do not introduce any data or models that can be deployed directly.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All reused assets are properly credited with licenses and terms explicitly stated:
    • Director Codebase: Built on the open-source implementation from [14], with modifications documented in the paper (Mentioned in the paper).
    • Miniworld Env: Apache 2.0 license
    No proprietary datasets or models were used.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We will provide the code as an asset with a comprehensive README.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: We do not use an LLM to develop any part of the method.